The linear regression equation for a data set is ˆy = 1.2 − 3.2 x. The actual value is -12.6, when  x = 5. What is the residual value at x=5?


residual =


1
Expert's answer
2021-11-23T16:12:46-0500

Given Regression Equation:

"y=1.2-3.2x"

Calculating the Predicted value at x =5 :

"y = 1.2 -3.2 \times 5 \\

y = -14.8"

The Actual Value at x = 5 is given : -12.6

residual = actual – predicted

residual "= -12.6 -(-14.8) = 2.2"

The Required Residual is 2.2